.Wednesday 31 July 2013: the start of the third week
On site: Keith Fitzpatrick-Matthews, Jim Skipper, Andrew Rylah, Christl Squires, Frankie Saxton, Ivor Davies, Jan Turner (until 1 pm), Jean Andrews, John Baskerville, Kit Carstairs, Paul Browne, Phil Dean, Steve Warner, Thomas Burningham, Tony Driscoll
Weather: overcast, frequent light drizzle, occasional light breeze; sunny spells from 11 o’clock and dry after 2.15, with the breeze becoming a little stiffer and more constant during the afternoon.
The weather today is quite unlike any we have had so far this year: it’s cool and damp, which is great for seeing the colour differences on site but not so pleasant for people working. At least it will clear during the day, although it is supposed to get very humid later.
Frankie is beginning to plan the north-western half of the site, so we ought to have a complete hand-drawn site plan in the next few days. With the colour differences being so much clearer today, it ought to be a relatively simple task.
Kit and Tom are working on the southern arm of the L-shaped section, removing a last little patch of topsoil (199) before tackling (197), the reddish deposit in the top of the inner ditch (described on last year’s sketch plan as “light grey”). In the damper conditions today, it is clear that (197) overlies (196) and that (97) is the material through which the inner ditch has been cut in this area. It is also likely that (197) is the same material as (195) to the north-east of the excavated section of inner ditch.
Ivor, Jean and Steve are working on what remains of (119). Fortunately, it seems to be quite superficial and is coming down rapidly onto (207), a much stonier deposit underneath, that also spills out through the gap in the bank.
Phil, Andrew and Paul are working on the outer ditch, . On Sunday, progress on the south-eastern part of (291) was much more rapid than to the north-west, leaving a step almost 0.25 m high between where different teams had been working. It is today’s priority task to get everything back to the same level, otherwise we risk exposing new deposits in one half of the ditch much earlier than in the other, which is quite unacceptable.
By the end of morning tea break, the drizzle had come to an end and temperature shot up by two degrees to 20° C. The sun is even coming through breaks in the cloud, so we may see an improvement in the weather rather earlier than the forecasts have been suggesting.
It seems to be quite a slow day for finds: there have been a few in (119) and (291), but (199) was without any. Kit and Tom have now removed the last patch of (119) and are on to (197), which does have material in it. We need to collect 20 litres of bulk sample from this deposit, as it is a fill of the inner ditch. The section excavated since 2011 has been assigned context ; we do not yet have a cut number for this section.
It’s fortunate that (291) has fewer finds as it becomes deeper, which means that progress can continue to be rapid in the outer ditch. We need to keep up our average of 4 cm a day to get to the bottom: should we lose any time to bad weather or extreme heat, we will need to increase the daily target. Comparing what we have this season with the section excavated through the outer ditch at Easter 2013, it is clear that (291) and (290) are identical with (285), which was removed entirely be machine; it was almost a metre thick at its deepest part.
The rain returned at lunchtime and was quite heavy for about forty minutes. By the time we were back at work, it had lightened and almost stopped, finally stopping again around 2.20 pm. The temperature has risen now to just touching 23° C; it drops as the wind increases, so we won’t be invoking the hot weather policy just yet.
The inner edge of the outer ditch is now showing up where (291) is being removed and is parallel with the opposite side, as one would expect. This shows that the point previously chosen as the mid point from which the two teams have been working back to the edge is too far towards the inner edge. Now that we know where the edges are, we will be able to set a more equitable starting point.
The rain over lunchtime has helped to show up the colours once again, helping to redefine areas where (119) has not yet been completely removed and to show up the loose rubbly material of (207) beneath it. The relationship between (207) and the rubbly deposits to its west is unclear: they appear to have less soil matrix, so I am inclined to regard them as something quite different.
There have been a number of sherds of shelly pottery again today, all coming from securely Neolithic deposits (both (197) and (119)). I wonder if they are made from the same Jurassic clays as the Romano-British shelly wares of North Bedfordshire and Northamptonshire; if this is the case, presumably they are part of a regional trade in ceramics. This is something I need to investigate further, as I am not familiar with this idea outside south-west England, where Gabbroitic Ware is “traded” over a wide area (scare quotes as we have no idea of the mechanisms by which goods moved from areas of production to areas of consumption).
One additional area we need to investigate is an extension to the sondage over the outer ditch. I would like to extend the edges to cross the remains of the bank, examine the relict topsoil beneath it and determine its relationship with the colluvium (35). This can then be added to the section drawings that will be made of ditch  when its excavation is complete, tying the outer ditch in stratigraphically to the henge proper.
Kit and Tom feel that they have found a new deposit under (197). Although the same colour, it has a higher proportion of chalk flecks and is somewhat firmer to trowel. Given the complex stratigraphy found in the section dug 2011-13, , it is unsurprising that they should find a new fill after only a little work.
Despite the rain, which was only really heavy at lunchtime, when we were sheltered under the gazebos, and the relative lack of people today, we seem to have made good progress today. The ‘step’ left in the outer ditch section where the two teams had been working at vastly different speeds on Sunday has now gone and we have the edge of the cut on its inner side. All of (119) has been removed, revealing that (207) lies directly beneath it, while the last remaining patch of topsoil (199) in the L-shaped section has also gone, enabling the team in this area to concentrate on the archaeological deposits of the inner ditch. It would be really good if, over the next four weeks, we can complete another section through the inner ditch, which would give us two sections through both ditches. We may have to make some compromises in how we plan each deposit to enable this to happen, for instance.
Posted on 31 July 2013, in Fieldwork, Stapleton's Field Dig 2013. Bookmark the permalink. Leave a comment.
Leave a comment